OpenStudy (anonymous):

Solve for the variable. -22 = x - 7 A. x = -15 B. x = -29 C. x = 15 D. x = 29

OpenStudy (anonymous):

Collect like terms

undeadknight26 (undeadknight26):

. -22 = x - 7 +7 + 7 ??? = x?

OpenStudy (anonymous):

\[-22 = x -7 \rightarrow x = -22 + 7 = -15\]

undeadknight26 (undeadknight26):

yup ^^^